# hide_content

### v2

@@ Abstract
There has been a surge of interest in language model agents that can navigate virtual environments such as the web or desktop. To navigate such environments, agents benefit from information on the various elements (e.g., buttons, text, or images) present. It remains unclear which element attributes have the greatest impact on agent performance, especially in environments that only provide a graphical representation (i.e., pixels). Here we find that the ordering in which elements are presented to the language model is surprisingly impactful—randomizing element ordering in a webpage degrades agent performance comparably to removing all visible text from an agent’s state representation. While a webpage provides a hierarchical ordering of elements, there is no such ordering when parsing elements directly from pixels. Moreover, as tasks become more challenging and models more sophisticated, our experiments suggest that the impact of ordering increases. Finding an effective ordering is non-trivial. We investigate the impact of various element ordering methods in web and desktop environments. We find that dimensionality reduction provides a viable ordering for pixel-only environments. We train a UI element detection model to derive elements from pixels and apply our findings to an agent benchmark—OmniACT—where we only have access to pixels. Our method completes more than two times as many tasks on average relative to the previous state-of-the-art.
@@ 1 Introduction
There has been growing interest in using language model (LM) agents to autonomously navigate virtual environments. Autonomous web agents [32, 10, 31, 5, 6] have become a particularly popular area of research. Typically, a web agent takes as input a task prompt from a user, observes a text and visual representation of the environment, and then outputs one or more actions to execute the task in the environment. Recently, research interest has expanded to include agents that can navigate mobile [15, 27] and desktop [26, 9] environments as well.
At a high level, a virtual environment consists of numerous elements—some are interactive (e.g. buttons or widgets), while others are not (e.g. plain text). To allow for human navigation, these elements are usually represented in the pixel space via a Graphical User Interface (GUI). In contrast, an agent must rely on some state representation to navigate a virtual environment. The exact format of a state representation varies between environments and approaches. In web environments, common text representations include the HTML or accessibility tree [32, 12]. For visual representations, a popular approach is to label UI elements with bounding boxes and numeric identifiers [28, 12, 6]. In either case, the state representation is derived from the underlying Document Object Model (DOM) [32, 12, 6], known as Set-of-Mark. However, constructing Set-of-Mark relies on knowledge of element locations, which are unknown by default in environments that only provide pixel information. To construct an effective state representation from only pixels, it is important to answer the following basic questions about these representations: What aspects of a state representation are most important to an agent? How can we derive these important aspects with only pixels?
@@
In almost all implementations of the agent’s state representation, there exists a list of interactable or non-interactable elements which the agent uses to determine the next action [12, 32, 6, 4, 27, 8, 9, 26]. Elements are characterized by various attributes such as visual appearance, text descriptions, or type labels. Because the state representation is the input to an LM, this list of elements is always presented in a specific ordering. For example, the default method to derive elements from a webpage performs a pre-order traversal of the Document Object Model (DOM) tree [23]. We analyze various attributes of a popular state representation for agents and find element ordering to be the single most impactful attribute to agent performance. We find that the ordering of elements can dramatically affect the performance of an agent, resulting in differences of up to 49% relative performance.
@@
The exact format of a state representation varies between environments and approaches. Common choices for a text representation include the HTML or accessibility tree [32, 12, 26]. For visual representations, a popular approach is to label UI elements with bounding boxes and numeric identifiers [28, 12, 6]. In almost all implementations, the agent observes a list of interactable or non-interactable elements to determine the next action. Elements are characterized by various attributes such as visual appearance, text descriptions, or type labels. Because the state representation is the input to an LM, this list of elements is always presented in a specific ordering. For example, the default method to derive elements from a webpage performs a pre-order traversal of the Document Object Model (DOM) tree [23]. We analyze various attributes of a popular state representation for agents and find element ordering to be the single most impactful attribute. . We find that the ordering of elements can dramatically affect the performance of an agent, resulting in differences of up to 49% relative performance.
This can prove problematic as many environments lack obvious methods to both derive and order elements. For example, many mobile applications do not properly expose interactable elements [2, 17, 30]. In such environments, pixels may be the only source of information available. Previous approaches to deriving interactable elements from pixels either leverage off-the-shelf segmentation models [27, 9] or build custom models that target accessibility features [24]. In our approach, we leverage common crawl [3] to train an object detection model [16] that detects interactable UI elements specifically for agents. To the best of our knowledge, the elements detected through these approaches are ordered arbitrarily (e.g. based on confidence scores); visually, the ordering is effectively random. Our experiments indicate that a random ordering consistently results in the lowest performance across multiple scenarios.
When not explicitly given, an effective ordering of elements is non-trivial to determine due to the hierarchical nature of most environments. Across multiple agent benchmarks, we find that ordering elements via a 2D-to-1D dimensionality reduction [21] reliably yields improvements to agent performance relative to other baselines. We experiment on the VisualWebArena [12] and OmniACT [9] benchmarks and achieve new state-of-the-art performance on OmniACT.
@@ 2 Related Work
World-of-bits provided the first environment for evaluating web GUI navigation using an agent [18]. Over time, more realistic web [32, 12, 29], desktop [9, 26], and mobile [15] agent benchmarks have been created. Kim et al. [10] provided one of the first LM agent approaches, successfully navigating World-of-bits. However, existing agents are still unable to properly navigate more realistic benchmarks, completing only 15% of web tasks [32] and 12% of desktop tasks.
@@
Despite work on multimodal agents [12], existing techniques in navigating web and desktop environments still rely heavily on ground-truth text representations. Zhou et al. [32], Koh et al. [12] both utilize the accessibility tree and its elements as their state representation. Koh et al. [12], He et al. [6] consider approaches where interactable elements are also represented in an image via Set-of-Mark [28] bounding boxes and labels. Xie et al. [26] provides an agent that navigates desktop applications by observing a filtered down version of the accessibility tree. All of these approaches require access to either a webpage’s underlying DOM or an accessibility tree to derive elements; however, our focus is on environments that only give access to their graphical representations which is significantly more challenging.
@@
There have been several approaches—primarily focused on desktop and mobile environments—to directly navigating a GUI via its pixels. Kapoor et al. [9] and Yan et al. [27] focus on navigating desktop and mobile applications respectively. Both leverage an off-the-shelf-segmentation model—Segment Anything [11]—to find icons in the image. These icons are then either represented in text [9] or labeled with Set-of-Mark [28] bounding boxes and labels in the image [27]. We instead train an object detection model that detects interactable UI elements directly. While previous UI element detection models are trained to detect accessibility features [24], our model is trained specifically to detect interactable elements that would be useful to an agent. All three of our approaches use Optical Character Recognition (OCR) modules such as EasyOCR [1] to extract text from pixel information.
@@
While most agent studies include some ablations, few focus on detailed analysis of an agent’s input. To our knowledge, Huq et al. [7] is the only other study that directly studies this. Their study focuses on broader components to an input prompt such as the selection of few-shot examples used, while we focus on specific element attributes such as element ordering.
@@ 3 Problem Definition
Elements can be represented in both visual and text modalities. For images, the most common approach is to overlay bounding boxes with numeric identifiers around each interactable element [12, 6, 27] in a manner inspired by Set-of-Mark Prompting [28]. In text, a common approach is to represent each element as a string containing its index, coordinates, and other attributes, such as "[1] [x,y] [Description]". Because LMs operate on sequential data, elements must be given an ordering; in most approaches, this is implicitly defined by the method used to identify the elements.
@@ 4 Which Aspects Of Agent State Representations Are Most Impactful?
Here we describe a series of ablation experiments designed to examine which aspects of an LM state representation are most impactful to the performance of LM agents. In particular, we experiment on the VisualWebArena (VWA) [12] benchmark and ablate attributes of the state representation of the state-of-the-art agent (proposed in the same paper). We pick this representation in particular due to both its strong performance on VWA, as well as its similarity to common practices seen in other agent research [6] and open source projects [8, 4]. Our experiments in turn ablate the impact of (1) multimodal (image and text) aspects of the state representation, and (2) individual element attributes within the text component alone.
@@ 4.1 VisualWebArena
VisualWebArena focuses on multimodal tasks in the web and provides a self-hostable environment for language agents to navigate [12]. Agents operating on VisualWebArena have full access to the DOM. The current best approach [12] utilizes a multimodal representation where elements are parsed in a pre-order traversal of the DOM tree [23]. Each element $j$ has attributes $S_{j}=\{id,tag,text\}$ where $id$ is a unique numeric identifier for interactable elements and $\varnothing$ otherwise, $tag$ is the HTML tag (e.g. BTN or IMG), and $text$ is the alt text and captions for images and the HTML text otherwise. In the text representation, an example of an element would be “[1] [IMG] [alt text, caption]”. In the image representation, each element is labeled with bounding boxes and numeric labels.
The original agent in [12] only achieves 15% success rate across all tasks. Since our goal is not to improve agent performance on VisualWebArena but rather to understand the importance of attributes in the state representation, we examine a subset of tasks to reduce costs. [^0] A full run of the state-of-the-art agent on VisualWebArena can cost up to \$800 with GPT-4V. Specifically, we explore tasks marked as “easy” within tasks that the original agent completed successfully. Due to variance associated with stochastic LM outputs, our reproduction of these originally-successful tasks yields a success rate around $74.07\%\pm 5.56\%$. The exact list of tasks can be found in Appendix A.4. We reuse the action space from the original agent which consists of executing high-level actions (e.g., click, hover) on individual elements—see Appendix A.2 for more details.
@@ 4.2 Ablation Setup and Findings
While adding a visual representation clearly improves performance, we find that it alone is insufficient even with Set-of-Mark labels. This contradicts previous findings on agents for mobile applications which found that a screenshot with Set-of-Mark labels achieves similar performance with or without text [27]. We speculate that this is due to the substantial difference in viewport sizes between mobile and desktop environments. Specifically, the average mobile device has a viewport size of 360x800 while the average desktop has a viewport size of 1920x1080 [20]. Additionally, larger viewport sizes have been shown to improve agent performance in desktop environments [26]. We speculate that this may be because current agents almost never understand when to change the screen view (e.g. by scrolling).
@@
Table 2: Ablating attributes of the text component of the VisualWebArena state representation. All results include the screenshot with Set-of-Mark bounding boxes and labels. TAG is the HTML tag. CAPTIONS are image captions generated using BLIP-2-T5XL[13]. TEXTAlt, TEXTInteract, and TEXTStatic are the alt text, text for interactable elements, and text for non-interactable elements respectively. ORDER is element ordering.  indicates ground truth obtained from the HTML. x indicates removal of the attribute. x Element Ordering indicates a random shuffling of the elements. - denotes experiments that were not run due to cost.
@@ 5 Experimental Setup
Most applications are built on top of an underlying hierarchical representation. For example, a webpage is modeled by the DOM which is hierarchical. When exposed, this hierarchy can be used to determine a strong element ordering. However, the availability and quality of an underlying hierarchical representation can vary greatly between environments. For example, Chen et al. [2] found that 77% of mobile applications had missing labels and Ross et al. [17] found that 53% of image buttons had missing labels and were incorrectly sized. Additionally, Ross et al. [17] found that $8\%$ of applications lacked interactable element information altogether. In such scenarios, we may only have access to the application’s pixel information. We continue to experiment with VisualWebArena and also experiment with the OmniACT benchmark as a scenario where we only have access to pixel information. Details on the LM agent backbones used in our experiments can be found in Appendix A.3.
@@ 5.1 OmniACT
To detect UI elements $\{e_{1},e_{2},\ldots,e_{n}\}$ when given only pixel information $\mathcal{P}$ we train an object detection model [16] to detect interactable UI elements in the screenshot and use EasyOCR [1] to extract visible text. In other words, the function $g\colon\mathcal{P}\rightarrow\mathcal{E}$ is defined by the trained object detection model. We add visible text and captions to each interactable UI element. We gather a dataset by finding 67,530 interactable elements over 1468 Common Crawl webpages. We selected our webpages based on top websites from Similarweb [19]. Despite the domain shift from webpages to desktop applications, we found that our object detection model worked reasonably well on the OmniACT benchmark in the end-to-end agent setting. We publicly release this model along with our paper. Training details can be found in the Appendix A.1.

### v3

@@ Abstract
There has been a surge of interest in language model agents that can navigate virtual environments such as the web or desktop. To navigate such environments, agents benefit from information on the various elements (e.g., buttons, text, or images) present. However, it remains unclear which element attributes have the greatest impact on agent performance, especially in environments that only provide a graphical representation (i.e., pixels). Here we find that the ordering in which elements are presented to the language model is surprisingly impactful—randomizing element ordering in webpages compromises average agent performance to a degree comparable to removing all visible text from webpages. While web agents benefit from the semantic hierarchical ordering of elements available via the browser, agents that parse elements directly from pixels do not have access to any such ordering. Here we endeavor to derive effective orderings and investigate the impact of various element ordering methods in web and desktop environments. We find that dimensionality reduction provides a viable ordering for pixel-only environments. We train a UI element detection model to derive elements from pixels and apply our findings to an agent benchmark—OmniACT—where we only have access to pixels. Our method completes more than two times as many tasks on average relative to the previous state-of-the-art.
@@ 1 Introduction
There has been growing interest in using language model (LM) agents to autonomously navigate virtual environments. Autonomous web agents (Zhou et al., 2023; Kim et al., 2023; Zheng et al., 2024; Gur et al., 2024; He et al., 2024) have become a particularly popular area of research. Typically, a web agent takes as input a task prompt from a user, observes a text and visual representation of the environment, and then outputs one or more actions to execute the task in the environment. Recently, research interest has expanded to include agents that can navigate mobile (Rawles et al., 2023; Yan et al., 2023) and desktop (Xie et al., 2024; Kapoor et al., 2024; Bonatti et al., 2024) environments as well.
At a high level, a virtual environment consists of numerous elements—some are interactive (e.g. buttons or widgets), while others are not (e.g. plain text). To allow for human navigation, these elements are usually represented in the pixel space via a Graphical User Interface (GUI). In contrast, agents often rely on distinct state representations to navigate virtual environments. The exact format of a state representation varies between environments and approaches. In web environments, common text representations include the HTML or accessibility tree (Zhou et al., 2023; Koh et al., 2024). For visual representations, a popular approach is to label UI elements with bounding boxes and numeric identifiers (Koh et al., 2024; He et al., 2024), known as Set-of-Mark (Yang et al., 2023). In either case, the state representation is derived from the underlying Document Object Model (DOM) (Zhou et al., 2023; Koh et al., 2024; He et al., 2024). However, many environments lack a descriptive DOM and only provide pixel information, which we show is insufficient for existing agents (see Section 4.2). To construct an effective state representation from only pixels, it is important to answer the following basic questions about these representations: What aspects of a state representation are most important to an agent? How can we derive these important aspects with only pixels?
@@
In almost all implementations of the agent’s state representation, there exists a list of interactable or non-interactable elements which the agent uses to determine the next action (Koh et al., 2024; Zhou et al., 2023; He et al., 2024; Dupont, 2024; Yan et al., 2023; Ishan0102, 2024; Kapoor et al., 2024; Xie et al., 2024). Elements are characterized by various attributes such as visual appearance, text descriptions, or type labels. Because the state representation is the input to an LM, this list of elements is always presented in a specific ordering. For example, the default method to derive elements from a webpage performs a pre-order traversal of the DOM tree (World Wide Web Consortium, 2013). We analyze various attributes of a popular state representation for agents and find element ordering to be the single most impactful attribute to agent performance. We find that the ordering of elements can dramatically affect the performance of an agent, resulting in differences of up to 49% relative performance.
@@
The exact format of a state representation varies between environments and approaches. Common choices for a text representation include the HTML or accessibility tree (Zhou et al., 2023; Koh et al., 2024; Xie et al., 2024). For visual representations, a popular approach is to label UI elements with bounding boxes and numeric identifiers (Yang et al., 2023; Koh et al., 2024; He et al., 2024). In almost all implementations, the agent observes a list of interactable or non-interactable elements to determine the next action. Elements are characterized by various attributes such as visual appearance, text descriptions, or type labels. Because the state representation is the input to an LM, this list of elements is always presented in a specific ordering. For example, the default method to derive elements from a webpage performs a pre-order traversal of the Document Object Model (DOM) tree (World Wide Web Consortium, 2013). We analyze various attributes of a popular state representation for agents and find element ordering to be the single most impactful attribute. . We find that the ordering of elements can dramatically affect the performance of an agent, resulting in differences of up to 49% relative performance.
This can prove problematic as many environments lack obvious methods to both derive and order elements. For example, many mobile and desktop applications (see Figure 2) do not properly expose interactable elements (Chen et al., 2020; Ross et al., 2020; Zhang et al., 2021). In such environments, pixels may be the only source of information available. Previous approaches to deriving interactable elements from pixels either leverage off-the-shelf segmentation models (Yan et al., 2023; Kapoor et al., 2024) or build custom models that target accessibility features (Wu et al., 2023). In our approach, we leverage common crawl (Common Crawl, 2023) to train an object detection model (Ren et al., 2016) that detects interactable UI elements specifically for agents. To the best of our knowledge, the elements detected through these approaches are ordered arbitrarily (e.g. based on confidence scores); visually, the ordering is effectively random. Our experiments indicate that a random ordering consistently results in the lowest performance across multiple scenarios.
Figure 2: Many software applications lack informative accessibility trees or DOMs. The accessibility tree for a popular game development engine (Godot, left) contains only the exit, minimize, and full screen buttons. For a presentation slide (Google Slides, right), no interactive elements (e.g. title and subtitle boxes) are present in the DOM. Language model agents rely on this information to navigate applications, and agent performance can accordingly be compromised in scenarios where it is incomplete.
Here we propose and evaluate strategies for deriving effective element orderings in scenarios where a hierarchical ordering based on the GUI design is not explicitly provided by the environment. Across multiple agent benchmarks, we find that ordering elements via a 2D-to-1D dimensionality reduction (Van der Maaten & Hinton, 2008) reliably yields improvements to agent performance relative to other baselines. We experiment on the VisualWebArena (Koh et al., 2024) and OmniACT (Kapoor et al., 2024) benchmarks and achieve new state-of-the-art performance on OmniACT.
@@ 2 Related Work
World-of-bits provided the first environment for evaluating web GUI navigation using an agent (Shi et al., 2017). Over time, more realistic web (Zhou et al., 2023; Koh et al., 2024; Yao et al., 2023), desktop (Kapoor et al., 2024; Xie et al., 2024), and mobile (Rawles et al., 2023) agent benchmarks have been created. Kim et al. (2023) provided one of the first LM agent approaches, successfully navigating World-of-bits. However, existing agents are still unable to properly navigate more realistic benchmarks, completing only 15% of web tasks (Zhou et al., 2023) and 12% of desktop tasks.
@@
Despite work on multimodal agents (Koh et al., 2024), existing techniques in navigating web and desktop environments still rely heavily on ground-truth text representations. Zhou et al. (2023); Koh et al. (2024) both utilize the accessibility tree and its elements as their state representation. Koh et al. (2024); He et al. (2024) consider approaches where interactable elements are also represented in an image via Set-of-Mark (Yang et al., 2023) bounding boxes and labels. Xie et al. (2024) provides an agent that navigates desktop applications by observing a filtered down version of the accessibility tree. All of these approaches require access to either a webpage’s underlying DOM or an accessibility tree to derive elements; however, our focus is on environments that only give access to their graphical representations which is significantly more challenging.
@@
There have been several approaches—primarily focused on desktop and mobile environments—to directly navigating a GUI via its pixels. Kapoor et al. (2024) and Yan et al. (2023) focus on navigating desktop and mobile applications respectively. Both leverage an off-the-shelf-segmentation model—Segment Anything (Kirillov et al., 2023)—to find icons in the image. These icons are then either represented in text (Kapoor et al., 2024) or labeled with Set-of-Mark (Yang et al., 2023) bounding boxes and labels in the image (Yan et al., 2023). We instead train an object detection model that detects interactable UI elements directly. While previous UI element detection models are trained to detect accessibility features (Wu et al., 2023), our model is trained specifically to detect interactable elements that would be useful to an agent. All three of our approaches use Optical Character Recognition (OCR) modules such as EasyOCR (AI, 2020) to extract text from pixel information.
@@
While most agent studies include some ablations, few focus on detailed analysis of an agent’s input. To our knowledge, Huq et al. (2023) is the only other study that directly studies this. Their study focuses on broader components to an input prompt such as the selection of few-shot examples used, while we focus on specific element attributes such as element ordering.
@@ 3 Problem Definition
Elements can be represented in both visual and text modalities. For images, the most common approach is to overlay bounding boxes with numeric identifiers around each interactable element (Koh et al., 2024; He et al., 2024; Yan et al., 2023) in a manner inspired by Set-of-Mark Prompting (Yang et al., 2023). In text, a common approach is to represent each element as a string containing its index, coordinates, and other attributes, such as "[1] [x,y] [Description]". Because LMs operate on sequential data, elements must be given an ordering; in most approaches, this is implicitly defined by the method used to identify the elements.
@@ 4 Which Aspects Of Agent State Representations Are Most Impactful?
Here we describe a series of ablation experiments designed to examine which aspects of an LM state representation are most impactful to the performance of LM agents. In particular, we experiment on the VisualWebArena (VWA) (Koh et al., 2024) benchmark and ablate attributes of the state representation of the state-of-the-art agent (proposed in the same paper). We pick this representation in particular due to both its strong performance on VWA, as well as its similarity to common practices seen in other agent research (He et al., 2024) and open source projects (Ishan0102, 2024; Dupont, 2024). Our experiments in turn ablate the impact of (1) multimodal (image and text) aspects of the state representation, and (2) individual element attributes within the text component alone.
@@ 4.1 VisualWebArena
VisualWebArena focuses on multimodal tasks in the web and provides a self-hostable environment for language agents to navigate (Koh et al., 2024). Agents operating on VisualWebArena have full access to the DOM. The current best approach (Koh et al., 2024) utilizes a multimodal representation where elements are parsed in a pre-order traversal of the DOM tree (World Wide Web Consortium, 2013). Each element $j$ has attributes $S_{j}=\{id,tag,text\}$ where $id$ is a unique numeric identifier for interactable elements and $\varnothing$ otherwise, $tag$ is the HTML tag (e.g. BTN or IMG), and $text$ is the alt text and captions for images and the HTML text otherwise. In the text representation, an example of an element would be “[1] [IMG] [alt text, caption]”. In the image representation, each element is labeled with bounding boxes and numeric labels.
The original agent in (Koh et al., 2024) only achieves 15% success rate across all tasks. Since our goal is not to improve agent performance on VisualWebArena but rather to understand the importance of attributes in the state representation, we examine a subset of tasks to reduce costs. [^0] A full run of the state-of-the-art agent on VisualWebArena can cost up to \$800 with GPT-4V. Specifically, we explore tasks marked as “easy” within tasks that the original agent completed successfully. Due to variance associated with stochastic LM outputs, our reproduction of these originally-successful tasks yields a success rate around $74.07\%\pm 5.56\%$. The exact list of tasks can be found in Appendix A.4. We reuse the action space from the original agent which consists of executing high-level actions (e.g., click, hover) on individual elements—see Appendix A.2 for more details.
@@ 4.2 Ablation Setup and Findings
While adding a visual representation clearly improves performance, we find that it alone is insufficient even with Set-of-Mark labels. This contradicts previous findings on agents for mobile applications which found that a screenshot with Set-of-Mark labels achieves similar performance with or without text (Yan et al., 2023). We speculate that this is due to the substantial difference in viewport sizes between mobile and desktop environments. Specifically, the average mobile device has a viewport size of 360x800 while the average desktop has a viewport size of 1920x1080 (Statcounter Global Stats, 2024). Additionally, larger viewport sizes have been shown to improve agent performance in desktop environments (Xie et al., 2024). We speculate that this may be because current agents almost never understand when to change the screen view (e.g. by scrolling).
Table 2: Ablating attributes of the text component of the VisualWebArena state representation. All results include the screenshot with Set-of-Mark bounding boxes and labels. TAG is the HTML tag. CAPTIONS are image captions generated using BLIP-2-T5XL(Li et al., 2023). TEXTAlt, TEXTInteract, and TEXTStatic are the alt text, text for interactable elements, and text for non-interactable elements respectively. ORDER is element ordering.  indicates ground truth obtained from the HTML. x indicates removal of the attribute. x Element Ordering indicates a random shuffling of the elements. - denotes experiments that were not run due to cost.
@@ 5 Experimental Setup
Most applications are built on top of an underlying hierarchical representation. For example, a webpage is modeled by the DOM which is hierarchical. When exposed, this hierarchy can be used to determine a strong element ordering. However, the availability and quality of an underlying hierarchical representation can vary greatly between environments. For example, Chen et al. (2020) found that 77% of mobile applications had missing labels and Ross et al. (2020) found that 53% of image buttons had missing labels and were incorrectly sized. Additionally, Ross et al. (2020) found that $8\%$ of applications lacked interactable element information altogether. In such scenarios, we may only have access to the application’s pixel information. We continue to experiment with VisualWebArena and also experiment with the OmniACT benchmark as a scenario where we only have access to pixel information. Details on the LM agent backbones used in our experiments can be found in Appendix A.3.
@@ 5.1 OmniACT
To detect UI elements $\{e_{1},e_{2},\ldots,e_{n}\}$ when given only pixel information $\mathcal{P}$ we train an object detection model (Ren et al., 2016) to detect interactable UI elements in the screenshot and use EasyOCR (AI, 2020) to extract visible text. In other words, the function $g\colon\mathcal{P}\rightarrow\mathcal{E}$ is defined by the trained object detection model. We add visible text and captions to each interactable UI element. We gather a dataset by finding 67,530 interactable elements over 1468 Common Crawl webpages. We selected our webpages based on top websites from Similarweb (2024). Despite the domain shift from webpages to desktop applications, we found that our object detection model worked reasonably well on the OmniACT benchmark in the end-to-end agent setting. We publicly release this model along with our paper. Training details can be found in the Appendix A.1.